# Online Difficulty Filtering for Reasoning Oriented Reinforcement Learning

### v1

@@ 6.3 Analysis
Figure 2: Reward as a function of step (2(a)) and relative time (2(b)). The horizontal dashed line indicates the maximum reward achieved by plain GRPO, and the vertical dashed lines indicate when GRPO with each threshold surpasses the plain GRPO’s maximum reward.
Online difficulty filtering adapts to model capability by presenting progressively harder examples.
Figure 3: Perceived difficulty per batch curated through balanced online filtering. Defining “difficulty” as $1-p(\mathbf{x})$ , a higher difficulty value implies lower sample accuracy.
@@
This observation, with the results in Table 1, shows that offline difficulty filtering with external proxies can provide partially meaningful difficulty assessments while not being perfectly aligned to the training model’s capability, shown through marginal improvements in Table 1 compared to plain GRPO. However, the advantage of the balanced online difficulty filtering is still evident in better benchmark results and efficiency.
@@ 7 Conclusion
We propose an online curriculum learning framework for reasoning-oriented reinforcement learning (RORL) in large language models (LLMs). By dynamically filtering training examples based on real-time pass rates, our approach ensures that the model focuses on problems within its optimal learning range. Experimental results demonstrate that this method improves sample efficiency and final model performance, outperforming both non-curriculum and offline curriculum baselines. Our findings underscore the importance of adaptive adjustment of training difficulty, paving the way for more effective reinforcement learning strategies for reasoning models.
Appendix A Asynchronous Implementation of Online Difficulty Filtering
@@
We provide a detailed diagram depicting the practical implementation of the online difficulty filtering, especially with the asynchronous setting (Noukhovitch et al., 2025).
@@
Figure 4: Illustration of the rollout process in the proposed algorithm with online difficulty filtering. Each iteration begins by sorting the dataset based on the visit count $\mathrm{vc(\mathbf{x})}$ of each example $\mathbf{x}$. A batch of unvisited or least-visited prompts is selected, respecting a predefined concurrency limit $C_{\text{max}}$. The asynchronous function $f_{async}$ samples responses from the current policy and evaluates them using the accuracy reward $r_{acc}$. Prompts with a pass rate within the accepted range $[T_{\text{Low}},T_{\text{High}}]$ are added to the training batch. Once the batch $\mathcal{B}$ reaches the target size $N$, any remaining asynchronous jobs in $\mathcal{P}_{\text{active}}$ are canceled. The policy is then updated using the GRPO loss computed over the collected batch.
Appendix B Learnability in Soft Prompts
Assuming that $r(\mathbf{x},\mathbf{y})\in\left\{0,1\right\}$ given the prompt $\mathbf{x}$ follows a Bernoulli distribution, we have:
$P(r(\mathbf{x},\mathbf{y})=1)=p(\mathbf{x})~{}~{}\mathrm{and}~{}~{}P(r(\mathbf{x},\mathbf{y}=0)=1-p(\mathbf{x}).$ (17)
Defining the inner term $\exp\left(\frac{1}{\beta}r(\mathbf{x},\mathbf{y})\right)$ as the random variable $Y$,
$Y=\begin{cases}~{}1&\mathrm{if}~{}~{}r(\mathbf{x},\mathbf{y})=0,\\ ~{}\exp\left(\frac{1}{\beta}\right)&\mathrm{if}~{}~{}r(\mathbf{x},\mathbf{y})=1.\end{cases}$ (18)
Thus, the expectation of $Y$ becomes:
$\mathbb{E}_{\mathbf{y}\sim\pi_{\mathrm{init}}\left(\cdot|\mathbf{x}\right)}\left[Y\right]$ $=\mathbb{E}_{\mathbf{y}\sim\pi_{\mathrm{init}}\left(\cdot|\mathbf{x}\right)}\left[\exp\left(\frac{1}{\beta}r(\mathbf{x},\mathbf{y})\right)\right]$ (19)
$=1\cdot P(r(\mathbf{x},\mathbf{y})=0)+\exp\left(\frac{1}{\beta}\right)\cdot P(r(\mathbf{x},\mathbf{y})=1)$ (20)
$=\left(1-p(\mathbf{x})\right)+p\cdot\exp\left(\frac{1}{\beta}\right),$ (21)
which leads to Equation (13) when applied to $V^{*}(\mathbf{x})$:
$V^{*}(\mathbf{x})=\beta\log\left((1-p(\mathbf{x}))+p(\mathbf{x})\exp\left(\frac{1}{\beta}\right)\right).$ (22)
Recall that:
$\mathbb{E}_{\mathbf{y}\sim\pi_{\mathrm{init}}(\cdot|\mathbf{x})}\left[\log\frac{\pi^{*}(\mathbf{y}|\mathbf{x})}{\pi_{\mathrm{init}}(\mathbf{y}|\mathbf{x})}\right]=\frac{1}{\beta}\mathbb{E}_{\mathbf{y}\sim\pi_{\mathrm{init}}(\cdot|\mathbf{x})}\left[r(\mathbf{x},\mathbf{y})\right]-\frac{1}{\beta}V^{*}(\mathbf{x}),$ (23)
we can substitute $\mathbb{E}_{\mathbf{y}\sim\pi_{\mathrm{init}}(\cdot|\mathbf{x})}\left[r(\mathbf{x},\mathbf{y})\right]$ with $p\left(\mathbf{x}\right)$ and $V^{*}(\mathbf{x})$ with Equation (22),
$\mathbb{E}_{\mathbf{y}\sim\pi_{\mathrm{init}}(\cdot|\mathbf{x})}\left[\log\frac{\pi^{*}(\mathbf{y}|\mathbf{x})}{\pi_{\mathrm{init}}(\mathbf{y}|\mathbf{x})}\right]=\frac{p(\mathbf{x})}{\beta}-\log\left((1-p(\mathbf{x}))+p(\mathbf{x})\exp\left(\frac{1}{\beta}\right)\right).$ (24)
For small $\frac{1}{\beta}$, the Taylor series expansion for the logarithm leads to:
$\log\left((1-p(\mathbf{x}))+p(\mathbf{x})\exp\left(\frac{1}{\beta}\right)\right)$ $=\log\left((1-p(\mathbf{x}))+p(\mathbf{x})\left(1+\frac{1}{\beta}+\frac{1}{2\beta^{2}}+\ldots\right)\right)$ (25)
$=\log\left(1+p(\mathbf{x})\left(\frac{1}{\beta}+\frac{1}{2\beta^{2}}+\ldots\right)\right)$ (26)
$\geq\log\left(1+p(\mathbf{x})\left(\frac{1}{\beta}+\frac{1}{2\beta^{2}}\right)\right).$ (27)
Since $\log\left(1+\epsilon\right)\geq\epsilon-\frac{\epsilon^{2}}{2}$, we can set $\epsilon=p(\mathbf{x})\left(\frac{1}{\beta}+\frac{1}{2\beta^{2}}\right)$:
$\log\left((1-p(\mathbf{x}))+p(\mathbf{x})\exp\left(\frac{1}{\beta}\right)\right)$ $\geq p(\mathbf{x})\left(\frac{1}{\beta}+\frac{1}{2\beta^{2}}\right)-\frac{1}{2}\left[p(\mathbf{x})\left(\frac{1}{\beta}+\frac{1}{2\beta^{2}}\right)\right]^{2}$ (28)
$=\frac{p(\mathbf{x})}{\beta}+\frac{p(\mathbf{x})}{2\beta^{2}}-\frac{p(\mathbf{x})^{2}}{2\beta^{2}}+\mathcal{O}\left(\frac{1}{\beta^{3}}\right)$ (29)
$=\frac{p(\mathbf{x})}{\beta}+\frac{p(\mathbf{x})(1-p(\mathbf{x}))}{2\beta^{2}}+\mathcal{O}\left(\frac{1}{\beta^{3}}\right).$ (30)
Substituting this back into the earlier equation, we obtain:
$\mathbb{E}_{\mathbf{y}\sim\pi_{\mathrm{init}}(\cdot|\mathbf{x})}\left[\log\frac{\pi^{*}(\mathbf{y}|\mathbf{x})}{\pi_{\mathrm{init}}(\mathbf{y}|\mathbf{x})}\right]$ $\leq\frac{p(\mathbf{x})}{\beta}-\left(\frac{p(\mathbf{x})}{\beta}+\frac{p(\mathbf{x})(1-p(\mathbf{x}))}{2\beta^{2}}\right)$ (31)
$=-\frac{p(\mathbf{x})(1-p(\mathbf{x}))}{2\beta^{2}}.$ (32)
Finally, recalling the definition of KL divergence,
$\mathbb{D}_{\mathrm{KL}}\left(\pi_{\mathrm{init}}(\mathbf{y}|\mathbf{x})\|\pi^{*}(\mathbf{y}|\mathbf{x})\right)=-\mathbb{E}_{\mathbf{y}\sim\pi_{\mathrm{init}}(\cdot|\mathbf{x})}\left[\log\frac{\pi^{*}(\mathbf{y}|\mathbf{x})}{\pi_{\mathrm{init}}(\mathbf{y}|\mathbf{x})}\right],$ (33)
We conclude that:

### v2

@@ 6.1 Learning dynamics analysis
Figure 3: Validation reward as a function of step (3(a)) and relative time (3(b)). The horizontal dashed line indicates the maximum reward achieved by plain GRPO, and the vertical dashed lines indicate when GRPO with each threshold surpasses the plain GRPO’s maximum reward.
@@
This observation, with the results in Table 1, shows that offline difficulty filtering with external proxies can provide partially meaningful difficulty assessments while not being perfectly aligned to the training model’s capability, shown through marginal improvements in Table 1 compared to plain GRPO. However, the advantage of the balanced online difficulty filtering is still evident in better performance and efficiency.
6.2 Training efficiency analysis
Figure 3 illustrates the progression of the reward in the validation set, plotted against both the training steps (3(a)) and the training time on the wall clock (3(b)). As shown in Figure 3(a), models trained with balanced online difficulty filtering consistently outperform the plain GRPO ($0\leq p(\mathbf{x})\leq 1$) in fewer training steps. For instance, using $0.4<p(\mathbf{x})<0.6$ achieved the highest overall reward and took the fewest gradient updates to outperform the plain GRPO within less than $50\%$ of steps in Figure 3(a). This suggests that by filtering out less informative examples, the average learnability within each batch increases, allowing faster learning progress. Interestingly, Figure 3(b) shows that this benefit carries over even when measured by wall-clock time by exceeding plain GRPO’s maximum reward in less training time. However, we also observe that overly aggressive filtering, such as in the case of the $0.4<p(\mathbf{x})<0.6$ setting, can require significantly more rollouts to fill a batch, leading to longer training times overall. These results suggest that online filtering can enable more efficient learning even in real-world settings.
6.3 Theoretical generalizability
We extend the learnability view of difficulty filtering in reinforcement learning for language models to a broader class of reward distributions. We begin by expressing the reverse KL divergence between the initial policy $\pi_{\mathrm{init}}(\cdot)$ and the optimal policy $\pi^{*}(\cdot)$ as the cumulant generating function (CGF) of a centered reward. We report the detailed proofs for the proposition and corollaries in Appendix G.
Proposition 6.1 (Reverse KL as CGF of centered reward).
Fix a prompt $x$ and temperature $\beta>0$. Let $\mu(x):=\mathbb{E}_{y\sim\pi_{\mathrm{init}}(\cdot|x)}[r(x,y)]$. Then
$\begin{split}&\mathbb{D}_{\mathrm{KL}}\!\big(\pi_{\mathrm{init}}(\cdot|x)\,\|\,\pi^{*}(\cdot|x)\big)\\ &=\log\,\mathbb{E}_{y\sim\pi_{\mathrm{init}}(\cdot|x)}\left[\exp\!\left(\frac{r(x,y)-\mu(x)}{\beta}\right)\right]\\ &=K_{\,r(x,\cdot)-\mu(x)}\!\left(\tfrac{1}{\beta}\right),\end{split}$ (17)
where $K_{Z}(t):=\log\mathbb{E}[\exp(tZ)]$ denotes the CGF of a random variable $Z$.
Proof sketch. Starting from the soft value $V^{*}(x)=\beta\log\mathbb{E}[e^{r(x,y)/\beta}]$ and the identity in equation 9,
$\begin{split}&\mathbb{D}_{\mathrm{KL}}\!\big(\pi_{\mathrm{init}}\|\pi^{*}\big)=-\mathbb{E}\!\left[\log\frac{\pi^{*}}{\pi_{\mathrm{init}}}\right]\\ &=\log\mathbb{E}\!\left[e^{r/\beta}\right]-\frac{1}{\beta}\,\mathbb{E}[r]\\ &=\log\mathbb{E}\!\left[e^{(r-\mu)/\beta}\right],\end{split}$ (18)
and the right-hand side is precisely the CGF of $r-\mu$ at $t=\tfrac{1}{\beta}$. $\square$
Proposition 6.1 highlights that learnability quantified by the reverse KL divergence is governed by the fluctuations of the reward around its mean. The variance term used as a proxy in §3 is the leading component of this control. This CGF view naturally generalizes to two practical settings: (1) Gaussian rewards, representing continuous rewards from neural evaluators (Lambert et al., 2025b; Wen et al., 2025); and (2) multinomial rewards, representing combinations of verifiable objectives (Guo et al., 2025; Gunjal et al., 2025).
Corollary 6.2 (Gaussian rewards).
If $r_{G}(x,y)\sim\mathcal{N}(\mu_{G}(x),\sigma_{G}^{2}(x))$ under $\pi_{\mathrm{init}}(\cdot|x)$, then
$\mathbb{D}_{\mathrm{KL}}\!\big(\pi_{\mathrm{init}}(\cdot|x)\,\|\,\pi^{*}(\cdot|x)\big)\;=\;\frac{\sigma_{G}^{2}(x)}{2\,\beta^{2}}.$ (19)
For Gaussian rewards, the reverse KL divergence, i.e., the learnability, is exactly proportional to the reward variance. Consequently, the sample variance from rollouts provides an unbiased and consistent empirical estimate of learnability. Let the sample variance
$\widehat{\sigma_{G}^{2}}(x)\;=\;\frac{1}{n-1}\sum_{i=1}^{n}\big(r_{i}-\bar{r}\big)^{2},$ (20)
where $r_{i}$ are $n$ independent reward samples. Since $\mathbb{E}[\widehat{\sigma_{G}^{2}}(x)]=\sigma_{G}^{2}(x)$,
$\mathbb{E}\!\left[\frac{1}{2\beta^{2}}\,\widehat{\sigma_{G}^{2}}(x)\right]=\mathbb{D}_{\mathrm{KL}}\!\big(\pi_{\mathrm{init}}(\cdot|x)\,\|\,\pi^{*}(\cdot|x)\big).$ (21)
Thus, the sample reward variance with sufficient size of rollouts can be a reliable proxy for learnability, even in the Gaussian reward setting, such as the classifier reward models (Lambert et al., 2025b).
Corollary 6.3 (Multinomial rewards).
Suppose $r_{M}(x,y)\in\{0,1,\dots,N\}$ with variance $\sigma^{2}(x)$ and cumulants $\kappa_{k}(x)$. Then, for $t=\tfrac{1}{\beta}$,
$\begin{split}&\mathbb{D}_{\mathrm{KL}}\!\big(\pi_{\mathrm{init}}(\cdot|x)\,\|\,\pi^{*}(\cdot|x)\big)=K_{\,r_{M}-\mu}\!(t)\\ &=\frac{\sigma^{2}(x)}{2}\,t^{2}+\frac{\kappa_{3}(x)}{6}\,t^{3}+\frac{\kappa_{4}(x)}{24}\,t^{4}+\cdots\\ &\geq\frac{\sigma^{2}(x)}{2\,\beta^{2}}-\mathcal{O}\!\left(\frac{1}{\beta^{3}}\right),\end{split}$ (22)
and in particular the variance $\tfrac{\sigma^{2}(x)}{2\beta^{2}}$ provides a tight second-order lower control of the reverse KL.
Here, the leading variance term is the dominant contribution to learnability, while the higher cumulants $\kappa_{3},\kappa_{4},\dots$ account for skewness and tail behavior typical of multi-objective or rubric-based rewards. The binary case is a special instance with N=1N{=}1. If $r_{B}(x,y)\in\{0,1\}$, then
$\mathrm{Var}[r_{B}(x,y)]\;=\;p(x)\big(1-p(x)\big),$ (23)
and substituting into the series above recovers the bound used in §3:
$\mathbb{D}_{\mathrm{KL}}\!\big(\pi_{\mathrm{init}}(\cdot|x)\,\|\,\pi^{*}(\cdot|x)\big)\;\geq\;\frac{p(x)\big(1-p(x)\big)}{2\,\beta^{2}}.$ (24)
This expands the applicability of the learnability viewpoint for the online difficulty filtering to the real-world setting, where multiple verifiable rewards are engaged, e.g., format rewards and accuracy rewards in reasoning (Guo et al., 2025; Luo et al., 2025; Wei et al., 2025a).
We further extend the principled discussion on algorithm-agnostic generalizability of the balanced online difficulty filtering in Appendix H.
@@ 7 Conclusion
In this work, we established the theoretical foundations of online difficulty-aware filtering in reinforcement learning with verifiable rewards (RLVR). Our analysis showed that tasks with intermediate difficulty maximize the lower bound of policy improvement, providing a principled explanation for the effectiveness of difficulty-based data curation, which was heuristically adopted in previous works. Through extensive ablation studies across multiple reasoning benchmarks, we verified these theoretical insights and demonstrated consistent gains in both sample efficiency and performance. These results highlight the importance of theoretically grounded difficulty control in RLVR, offering a unified perspective that connects empirical heuristics with formal learning principles.
Limitations
Our work provides both theoretical and empirical guidelines for online difficulty filtering in reasoning-oriented reinforcement learning for language models. While our theoretical analysis can be applied to any verifiable task, as we have shown through the proposition and corollaries, our empirical validation was conducted solely on math reasoning tasks. We leave the exploration of diverse verifiable tasks, such as coding and scientific reasoning, for future work. Furthermore, we plan to investigate the broader applicability of our method to larger scales and wider model families.
Appendix A Related Works
Reinforcement learning with verifiable rewards.
Recent advancements demonstrate significant reasoning improvements in LLMs through RL (Havrilla et al., 2024; OpenAI et al., 2024; Lambert et al., 2024; Guo et al., 2025; Walsh et al., 2025; Kumar et al., 2025). OpenAI o1 (OpenAI et al., 2024) initially reported that increasing the compute during RL training and inference improves reasoning performance. DeepSeek R1 (Guo et al., 2025) further found that, in reinforcement learning with verifiable rewards (RLVR), longer responses correlate with better reasoning. Concurrent studies (Team et al., 2025; Hou et al., 2025; Luo et al., 2025) employed algorithms, such as GRPO (Shao et al., 2024) or RLOO (Ahmadian et al., 2024), relying on advantage estimation via sampling rather than PPO-like value networks. Hou et al. (2025) further found that training efficiency improved with increased sampling in RLOO, invoking the need for more sample-efficient training strategies in RLVR.
Difficulty-based curriculum learning.
Curriculum learning has been widely adopted in fine-tuning LLMs to improve training efficiency (Lee et al., 2024; Naïr et al., 2024; Team et al., 2025; Cui et al., 2025). Static curricula, i.e., offline data curation with a predetermined task difficulty, have been effective in multiple domains: instruction-tuning (Lee et al., 2024) and coding (Naïr et al., 2024; Team et al., 2025; Li et al., 2025) to name a few. In RLVR, Team et al. (2025) employs a static difficulty-based curriculum, assigning tasks at fixed difficulty levels to ensure efficient progression. Similarly, Li et al. (2025) selects a high-impact subset of training data based on a “learning impact measure”. Meantime, adaptive curricula dynamically adjust task difficulty based on the learners’ progress, addressing the limitations of static curricula (Florensa et al., 2018; Cui et al., 2025). Specifically, Cui et al. (2025) applied adaptive filtering in reasoning and reported an empirical advantage in reducing reward variance. However, Meng et al. (2025) observed that such dynamic exclusion of examples may destabilize training, as it causes fluctuations in the effective batch size.
These works introduce adaptive filtering heuristics for RLVR but do not characterize when and why such filtering helps or how it interacts with KL-regularized RL objectives. In contrast, our work (1) derives a reverse-KL lower bound that identifies reward variance as the key learnability proxy, (2) shows that this directly motivates intermediate-difficulty selection, and (3) proposes an asynchronous fixed-batch implementation that avoids the batch-size instability.
Appendix B Asynchronous Implementation of Online Difficulty Filtering
@@
We provide a detailed diagram depicting the practical implementation of the online difficulty filtering, especially with the asynchronous setting (Noukhovitch et al., 2025). The formal expression of filling the batch $\mathcal{B}^{(t)}$ for the balanced online difficulty filtering is:
Here, the sample mean of $r_{acc}(\mathbf{x},\mathbf{y}_{i})$ is an unbiased estimate of $\mathbb{E}_{\mathbf{y}\sim\pi_{\theta_{t}}(\cdot|\mathbf{x})}\left[r_{\mathrm{acc}}(\mathbf{x},\mathbf{y})\right]$. Rollouts for each prompt are sampled asynchronously and in parallel, enabling continuous batching of prompts and rollouts (Daniel et al., 2023; Kwon et al., 2023; Noukhovitch et al., 2025). Each prompt’s visit count, $\mathrm{vc}(\mathbf{x})$, is incremented after generating $G$ rollouts, ensuring it isn’t re-processed in the same iteration. Moreover, the active rollout process $\mathcal{P}_{\text{active}}$ is halted once the batch capacity is reached, allowing prompt training with the collected data. This sampling-based framework is compatible with Monte Carlo methods such as RLOO (Ahmadian et al., 2024) and VinePPO (Kazemnejad et al., 2024).
@@
Figure 4: Illustration of the rollout process in the proposed algorithm with online difficulty filtering. Each iteration begins by sorting the dataset based on the visit count $\mathrm{vc(\mathbf{x})}$ of each example $\mathbf{x}$. A batch of unvisited or least-visited prompts is selected, respecting a predefined concurrency limit $C_{\text{max}}$. The asynchronous function $f_{async}$ samples responses from the current policy and evaluates them using the accuracy reward $r_{acc}$. Prompts with a pass rate within the accepted range $T_{\text{Low}},T_{\text{High}}]$ are added to the training batch. Once the batch $\mathcal{B}$ reaches the target size NN, any remaining asynchronous jobs in $\mathcal{P}_{\text{active}}$ are canceled. The policy is then updated using the GRPO loss computed over the collected batch.
Appendix C Learnability in Soft Prompts
We prove the variance-controlled lower bound in Proposition 3.1.
Proof. Throughout, let $r_{\mathrm{acc}}(x,y)\in\{0,1\}$ denote a binary verifiable reward with
$P\!\big(r_{\mathrm{acc}}(x,y)=1\big)=p(x)~~\text{and}~~P\!\big(r_{\mathrm{acc}}(x,y)=0\big)=1-p(x),$ (26)
and assume $\beta>0$ with $1/\beta\ll 1$ so that second-order Taylor expansions are valid. Expectations are taken with respect to $y\sim\pi_{\mathrm{init}}(\cdot\mid x)$ unless noted.
Step 1: Soft value under Bernoulli rewards.
Define the random variable
$Y\;=\;\exp\!\Big(\tfrac{1}{\beta}\,r_{\mathrm{acc}}(x,y)\Big)\;=\;\begin{cases}1,&r_{\mathrm{acc}}(x,y)=0,\\ \exp(1/\beta),&r_{\mathrm{acc}}(x,y)=1.\end{cases}$ (27)
Then
$\mathbb{E}[Y]\;=\;(1-p(x))+p(x)\,e^{1/\beta}.$ (28)
Using the soft value definition, we obtain
$V^{*}(x)\;=\;\beta\,\log\Big((1-p(x))+p(x)\,e^{1/\beta}\Big).$ (29)
Step 2: Expected log-ratio.
From the identity in equation 9, substituting $\mathbb{E}[r_{\mathrm{acc}}(x,y)]=p(x)$ and equation 29 gives
$\mathbb{E}\!\left[\log\frac{\pi^{*}(y\mid x)}{\pi_{\mathrm{init}}(y\mid x)}\right]\;=\;\frac{p(x)}{\beta}-\log\!\Big((1-p(x))+p(x)\,e^{1/\beta}\Big).$ (30)
Step 3: Second-order expansion and lower bound.
Write $e^{1/\beta}=1+\tfrac{1}{\beta}+\tfrac{1}{2\beta^{2}}+\mathcal{O}(\beta^{-3})$. Then
$(1-p(x))+p(x)\,e^{1/\beta}\;=\;1+p(x)\!\left(\frac{1}{\beta}+\frac{1}{2\beta^{2}}\right)+\mathcal{O}\!\left(\frac{1}{\beta^{3}}\right).$ (31)
Using $\log(1+\epsilon)\geq\epsilon-\tfrac{\epsilon^{2}}{2}$ with $\epsilon=p(x)\!\left(\tfrac{1}{\beta}+\tfrac{1}{2\beta^{2}}\right)$ yields
$\log\!\Big((1-p(x))+p(x)\,e^{1/\beta}\Big)\;\geq\;\frac{p(x)}{\beta}+\frac{p(x)\big(1-p(x)\big)}{2\beta^{2}}+\mathcal{O}\!\left(\frac{1}{\beta^{3}}\right).$ (32)
Subtracting equation 32 from equation 30 gives